# A comprehensive survey on deep active learning and its applications in medical image analysis

### v1

@@ 4 Integration of Active Learning and Other Label-Efficient Techniques
Wu et al. (2022c) 2022 MICCAI - - Cover-based Cosine Similarity Diversity - Clustering Slice
Bai et al. (2022) 2022 MICCAI Model Disagreement Entropy-weighted Dice Distance - - Hybrid - Fusion Pseudo-label
@@
[…]
Lyu et al. (2023) 2023 CVPR Data Disagreement Cross Entropy, Variance - - Hybrid - Fusion Pseudo-label Box
@@
[…]
@@
Sun et al. (2023) 2023 ICLR Uncertainty-aware Model - EDL Evidential Uncertainty Density-based Inverse Average Distance Hybrid Fusion
Sadafi et al. (2023) 2023 ISBI Multiple Inferences - MC Dropout Model Disagreement Variance Inequality - - Hybrid - Fusion
[…]
@@ 4.1 Semi-supervised Learning: Utilizing Unlabeled Data
Semi-supervised learning (Chen et al., 2022a) aims to boost performance by utilizing unlabeled data upon supervised training. AL and semi-supervised learning complements each other: AL focuses on constructing an optimal labeled dataset. However, massive unlabeled samples are discarded during model training. Therefore, we can further leverage unlabeled data to train the deep model. By integrating the strengths of both AL and semi-supervised learning, annotation efficiency can be further improved. This section will introduce the integration of AL and semi-supervised learning from pseudo-labeling and consistency regularization.
@@ 4.1.1 Pseudo-Labeling
Pseudo-labeling (Lee et al., 2013) is one of the most straightforward methods in semi-supervised learning. It uses the model’s predictions of unlabeled data as pseudo-labels and combines them with labeled data for supervised training. Although it’s possible to assign pseudo-labels to all unlabeled samples for training, it may introduce noise. To mitigate this, Wang et al. (2017) proposed cost-effective active learning (CEAL), integrating pseudo-labeling with uncertainty-based AL. Specifically, CEAL sent the most uncertain samples for expert annotation and assigned pseudo-labels to the most confident samples. Many subsequent works have built upon the ideas of CEAL. In point cloud segmentation, both Hu et al. (2022) and Liu et al. (2022a) assigned pseudo-labels of the most certain regions. In medical image segmentation, Zhao et al. (2021) refined the pseudo-labels with dense conditional random fields. Additionally, Li et al. (2022) proposed a new approach for selecting samples for oracle annotation and pseudo-label. Specifically, they employed curriculum learning to categorize all samples into hard and easy. Hard samples were all sent for oracle annotation. For the easy samples, they evaluated the presence of label noise based on the training loss. Easy samples with low training loss were used for pseudo-labels to assist training, whereas easy samples with high loss were considered noisy and excluded from training.
@@ 4.1.2 Consistency Regularization
Consistency regularization is also widely applied in semi-supervised learning. Its basic idea is to enforce consistent outputs under perturbations of input data or model parameters. Maximizing consistency serves as an unsupervised loss for unlabeled samples. Consistency regularization helps improve the robustness and reduce overfitting of the model, thus enhancing model performance. Gao et al. (2020) introduced a semi-supervised active learning framework. Consistency here was used for both semi-supervised training and evaluating informativeness. In this framework, samples are fed into the model multiple times with random augmentations. The consistency loss of unlabeled samples was implemented by minimizing the variance between multiple outputs. They further selected less consistent samples for annotation. Results showed that combining AL with semi-supervised learning significantly improves performance.
Additionally, some works integrated existing consistency-based semi-supervised methods into the training process of AL. Huang et al. (2021) combined their proposed COD with MeanTeacher (Tarvainen and Valpola, 2017), demonstrating superior performance. Both TypiClust (Hacohen et al., 2022) and ProbCover Yehuda et al. (2022) found that their methods outperformed other active learning baselines in low-budget scenarios when combined with FlexMatch (Zhang et al., 2021). Wang et al. (2022c) combined density-based AL with different existing semi-supervised methods. Results showed that the proposed method outperforms other active learning methods and excels in semi-supervised learning. Zhang et al. (2022a) combined AL with both pseudo-labeling and consistency. The unlabelled images first underwent both strong and weak data augmentations. When the confidence level of the weakly augmented images exceeded a certain threshold, they used these samples for semi-supervised training. Specifically, predictions of the weakly augmented images were assigned as pseudo-labels, and the outputs of the strongly augmented images were forced to be consistent with the pseudo-labels. However, when the confidence level was lower than the threshold, they used these samples for AL. A balanced uncertainty selector and an adversarial instability selector were used to select samples for oracle annotation. They validated the effectiveness of their proposed method in grading metastatic epidural spinal cord compression with MRI images.
@@ 4.2 Self-supervised Learning: Utilizing Pre-trained Model
In the last section, we discussed the integration of AL and semi-supervised learning, which aimed to utilize unlabeled data for better performance. However, the effectiveness of this approach is constrained by the size of dataset. This limitation is particularly evident in medical image analysis, where datasets are often relatively small. To further improve annotation efficiency, AL can be combined with self-supervised learning. Self-supervised learning (Liu et al., 2021a) trained the model with the supervision from the data itself, thus allowing pre-training on a large dataset. After finetuning on a few randomly selected labeled samples, the self-supervised pre-trained models have been shown to achieve impressive performance (Chen et al., 2020). Besides, these models can also provide good initialization, thereby solving the cold-start problem in AL. In this section, we will first introduce how self-supervised models solve the cold-start problem in AL and then explore different ways of integrating active learning with self-supervised learning.
@@ 4.2.1 Cold-start Problem in Active Learning
Current AL methods usually require several initial labeled samples to train an initial model and ensure reliable informativeness metrics. However, when the initial labeled set is small or even absent, the performance of these AL methods drops dramatically, sometimes even worse than random sampling (Chen et al., 2023; Hacohen et al., 2022; Yehuda et al., 2022). This is known as the cold-start problem in AL, which is very common in AL. In their semi-supervised active learning framework, Gao et al. (2020) found that performance suffered with a smaller labeling budget compared to a larger one when initial labels were randomly selected. Bengar et al. (2021) first pre-trained the model with self-supervised learning and then employed some AL baselines to select samples for labeling and finetuning. Results showed the performance of AL baselines tends to be worse than the random selection under the low-budget scenario. Additionally, Xie et al. (2023a) discovered that CoreSet based on self-supervised features is inferior to random sampling. Tackling the cold-start problem is vital for improving the efficacy of AL, especially when the annotation budget is limited. Moreover, when constructing a new dataset from scratch, employing cold-start AL strategies can offer a good initialization, thereby boosting performance.
A key solution to the cold-start problem in AL lies in selecting the optimal set of initial labeled samples. Since no initial labels are available, cold-start AL requires different strategies than existing AL methods. Self-supervised pre-trained models offer a good initialization for effectively tackling the cold-start problem in AL. In natural language processing, Yuan et al. (2020) was the first to introduce the cold-start problem in AL. They employed self-supervised pre-trained models to address this issue. Yi et al. (2022) chose initial samples based on the loss of self-supervised pretext tasks, showing significant advantages over random sampling. Pourahmadi et al. (2021) proposed a straightforward baseline for cold-start active learning. They first performed k-Means clustering on existing off-the-shelf self-supervised features, then selected cluster centers for annotation. Results indicated that this baseline is very effective when the annotation budget is limited. TypiClust (Hacohen et al., 2022) found that when the annotation budget is low, querying typical samples is more beneficial, whereas when the budget is high, querying hard samples is more beneficial. This conclusion suggested different strategies with uncertainty methods for cold-start AL. Thus, based on self-supervised features, TypiClust selected samples from high-density areas of each k-Means cluster. Yehuda et al. (2022) employed a graph-based greedy algorithm to select the optimal initial samples based on self-supervised features. Chen et al. (2023) found that active learning also suffers from a cold-start problem in medical image analysis. The issues arose mainly because AL is often biased towards specific classes, resulting in class imbalance. Additionally, the models struggled to detect anomalies when only a limited number of initially labeled samples exist. They combined clustering and the loss of contrastive learning to address the cold-start problem. In CT segmentation, Nath et al. (2022) designed new pretext tasks for self-supervised pre-training. The model was trained to learn the threshold segmentation by an abdominal soft-tissue window. Results indicated that the proposed method significantly outperforms random sampling in selecting initial samples.
@@
Additionally, some works have attempted to use fully supervised pre-trained models to address the cold-start problem. Zhou et al. (2017) and their subsequent work (Zhou et al., 2021b) used ImageNet pre-trained models to select samples for labeling from completely unlabeled datasets. They combined entropy and disagreement as informativeness metrics, where the disagreement was the KL divergence of prediction probabilities between different patches of the same sample. They also introduced randomness to balance exploration and exploitation.
@@ 4.2.2 Combination of Active Learning and Self-supervised Learning
Features: The simplest way to integrate AL with self-supervised learning is by leveraging the high-quality pre-trained features that effectively capture data similarities. In point cloud segmentation, Hou et al. (2021) performed k-Means clustering on the self-supervised features, then selected the points of cluster centers for annotation. They improved the annotation efficiency in indoor scene point cloud segmentation.
@@
Pretext tasks: In addition, the pretext tasks in self-supervised learning can be used for AL. They are tasks for which the supervision comes directly from the data itself. Different pretext tasks correspond to different pre-training paradigms. Typical pretext tasks include rotation prediction (Gidaris et al., 2018), colorization (Zhang et al., 2016), jigsaw puzzles (Noroozi and Favaro, 2016), contrastive learning (He et al., 2020), and masked modeling (He et al., 2022), etc. Solving these pretext tasks on extensive unlabeled data, the model acquires useful feature representations that can indirectly reflect data characteristics. Related works generally employed the loss of pretext task for AL. Yi et al. (2022) found a strong correlation between the loss of pretext tasks and the loss of downstream tasks. Thus, they initially focused on annotating samples with higher loss of pretext tasks and later shifted to those with lower loss. Results showed that rotation prediction performed the best among different pretext tasks. In Chen et al. (2023), the loss of contrastive learning was used for AL. They assumed that samples with higher losses are more representative of the data distribution. Specifically, they pre-trained on the target dataset using MoCo (He et al., 2020) for contrastive learning and then used k-Means clustering to partition the unlabeled data into multiple clusters, selecting the samples with the highest contrastive loss within each cluster for annotation. They then selected samples with the highest contrastive loss in each cluster for annotation.
@@
Others: Furthermore, we can also leverage self-supervised learning in other ways for AL. In classification tasks, Zhang et al. (2022b) introduced one-bit annotation into AL for classification tasks. Firstly, they selected informative samples through uncertainty metrics. Oracles returned whether the current prediction was right or wrong rather than full annotation. Then, contrastive learning was adopted to pull the correct predictions closer to their corresponding classes and push away wrongly predicted samples from the predicted classes. Results indicated that the proposed method outperforms other AL methods regarding bit information. Du et al. (2021) integrated contrastive learning into AL to tackle the problem of class distribution mismatch, where unlabeled data often includes samples out of the class distribution of the labeled dataset. In this work, contrastive learning filtered samples of mismatched classes that differ from the current class distribution. Besides, contrastive learning highlighted samples’ informativeness by setting carefully designed negative samples. Their extended work Du et al. (2022) provided more theoretical analysis and experimental results and further integrated existing label information into the contrastive learning framework.
4.3 Active Domain Adaptation: Tackling Distribution Shift
Domain Adaptation (DA) (Guan and Liu, 2021) has wide applications in medical image analysis and computer vision. It aims to transfer knowledge from the source to the target domain, thus minimizing annotation costs. Currently, the most common setting of DA is unsupervised domain adaptation (UDA), in which the source domain is labeled while the target domain is unlabeled. However, the performance of UDA still lags behind fully supervised learning in the target domain (Liu et al., 2023). To bridge this gap, a natural idea would be to employ AL to select and annotate informative samples in the target domain. This setting is known as active domain adaptation (ADA). For better queries in ADA, one should consider uncertainty and whether the sample represents the target domain. The latter is commonly referred to as domainness or targetness in ADA. This section reviews the development of ADA and explores various ways of integrating AL with DA.
Su et al. (2020) was the first to introduce the concept of ADA and combined domain adversarial learning with AL. Through a domain discriminator and task model, they performed importance sampling to select target domain samples that are uncertain and highly different from the source domain. Fu et al. (2021) combined query-by-committee, uncertainty, and domainness for selecting the most informative samples under distribution shift. They adopted a domain discriminator to select samples with high domainness and employed Gaussian kernels to filter out anomalous and source-similar samples of the target domain. Random sampling was also used to improve diversity. Prabhu et al. (2021) performed k-Means clustering on target domain samples and selected cluster centers for annotation. The cluster centers were weighted by uncertainty, thus ensuring that selected samples were uncertain and diverse. Rangwani et al. (2021) formulated ADA as a submodular optimization problem. The sum of uncertainty, diversity, and representativeness was considered the gain for annotating a sample. Specifically, uncertainty was measured by the KL divergence between the original samples and their adversarial samples. Diversity was defined as the minimum KL divergence from a single sample to a set of samples. The Bhattacharya coefficient between samples was used as the representativeness score. They adopted a greedy algorithm to iteratively pick samples with the maximum gain. In segmentation tasks, Ning et al. (2021) introduced the idea of anchors in ADA. They concatenated features of different classes from the source domain images. Cluster centers of these concatenations were referred to as anchors. They then computed the distance between each target sample and its nearest anchor. Target samples with the highest distance were requested for annotation. Shin et al. (2021) proposed LabOR, which first used a UDA pre-trained model to generate pseudo-labels for target samples and trained two segmentation heads with these pseudo-labels. They maximized the disagreements between the two heads and annotated regions that exhibited the most disagreement. LabOR achieved performance close to full supervision with only 2.2% of target domain annotations. Hwang et al. (2022) first selected representative samples in the target domain with maximum mean discrepancy. Uncertain ones within these samples were sent for annotation, while the confident ones are used for pseudo-labels. Xie et al. (2022b) introduced the concept of energy (LeCun et al., 2006) into ADA. The energy is inversely proportional to the likelihood of the data distribution. In this work, the model trained on the source domain was used to calculate the energy of target domain samples. Samples with high energy were selected for annotation, which suggested they are representative of the target domain and substantially different from the source data. Xie et al. (2022d) spotted the hard source samples by maximizing margin loss and leveraged these samples to select target samples close to the decision boundary. Based on EDL, Xie et al. (2022c) incorporated Dirichlet distribution to mitigate model miscalibration on the target domain. Distribution and data uncertainty were both used for sample selection. Huang et al. (2023) selected samples with high uncertainty and prediction inconsistency to their nearest prototypes. In 3D object detection, Yuan et al. (2023) adopted a diversity-based strategy to select target domain samples. Specifically, they first clustered samples based on similarity and selected prototypes of each cluster for annotation.

### v3

@@ 4 Integration of Active Learning and Other Label-Efficient Techniques
Wu et al. (2022d) 2022 MICCAI - - Cover-based Cosine Similarity Diversity - Clustering Slice
Bai et al. (2022) 2022 MICCAI Multiple Inferences - Model Disagreement Entropy-weighted Dice Distance - - Hybrid - Fusion Pseudo-label
@@
Xie et al. (2023b) 2023 CVPR - - Discrepancy-based Wasserstein Distance Latent Space Optimization & Nearest Neighbour Search
Lyu et al. (2023) 2023 CVPR Multiple Inferences - Data Disagreement Cross Entropy, Variance - - Hybrid - Fusion Pseudo-label Box
@@
Xie et al. (2022c) 2023 ICLR Uncertainty-aware Model - EDL Mutual Information & Entropy Expectation of Dirichlet Distribution - - Hybrid - Multi-round
@@
Sadafi et al. (2023) 2023 ISBI Multiple Inferences - MC Dropout Multiple Inferences - Model Disagreement Variance Inequality - - Hybrid - Fusion
Bai et al. (2023) 2023 MICCAI Multiple Inferences - Model Disagreement Gradient-based Uncertainty KL divergence, gradient Cover-based L2 Distance Diversity - Clustering
Tang et al. (2023) 2023 MICCAI Multiple Inferences - Model Disagreement KL divergence - - Top-k
Qiu et al. (2023) 2023 MICCAI Single Model Distance to 0.5 - - Top-k Patch
Chen et al. (2023b) 2023 MIDL - - Loss of Self-supervised Pretext Tasks Top-k
Qu et al. (2023a) 2023 NeurIPS Multiple Inferences - Model Disagreement Variance, Entropy, Overlap - - Top-K
Lou et al. (2023) 2023 TMI - - Clustering-based Consistency Diversity - Clustering Pseudo-label
Du et al. (2022) 2023 TPAMI - - Discrepancy-based Semantic and distinctive scores Hybrid - Fusion
@@ 4.1 Semi-supervised Learning: Utilizing Unlabeled Data
Semi-supervised learning (Chen et al., 2022a; Han et al., 2024) aims to boost performance by utilizing unlabeled data upon supervised training. The need for tedious human annotation can be further reduced by integrating AL and semi-supervised learning in medical image analysis. The reason is that AL and semi-supervised learning complement each other. Specifically, a large pool of unlabeled images should be collected from the hospital information systems to train a DL model for some clinical applications. With the help of AL, the DL model is trained on an optimal labeled dataset constructed with a certain AL method, which reduces the annotation workload for doctors. However, massive unlabeled samples sit idle during the model training in the AL cycle. By combining the AL with semi-supervised learning, the model can be trained on both labeled and unlabeled samples. (Jiménez et al., 2023) This section will introduce the integration of AL and semi-supervised learning from the perspectives of pseudo-labeling and consistency regularization.
@@ 4.1.1 Pseudo-Labeling
Pseudo-labeling (Lee et al., 2013) is one of the most straightforward methods in semi-supervised learning. It uses the model’s predictions of unlabeled data as pseudo-labels and combines them with labeled data for supervised training. Although it’s possible to assign pseudo-labels to all unlabeled samples for training, it could introduce noise. To mitigate this, Wang et al. (2017) proposed cost-effective active learning (CEAL), integrating pseudo-labeling with uncertainty-based AL. Specifically, CEAL sent the most uncertain samples for expert annotation and assigned pseudo-labels to the most confident samples. Many subsequent works have built upon the ideas of CEAL. Gorriz et al. (2017) adopted the CEAL framework in the melanoma segmentation and used the MC dropout for uncertainty estimation. In medical image segmentation, Zhao et al. (2021) refined the pseudo-labels with dense conditional random fields. Additionally, Li et al. (2022) proposed a new approach for selecting samples for oracle annotation and pseudo-labeling in the Gleason grading of prostate cancer with histopathology images. They employed curriculum learning to categorize all samples into hard and easy. Hard samples were all sent for oracle annotation. For the easy samples, they evaluated the presence of label noise based on the training loss. Easy samples with low training loss were used for pseudo-labels to assist training, whereas easy samples with high loss were considered noisy and excluded from training.
@@ 4.1.2 Consistency Regularization
Consistency regularization aims to enforce similar outputs under perturbations of input data or model parameters. Maximizing consistency serves as an unsupervised loss for unlabeled samples, which helps improve the robustness, reduces overfitting, and improves model performance. Many works integrated existing consistency-based semi-supervised methods into the training process of AL. In chest X-ray classification, Balaram et al. (2022) incorporated several semi-supervised methods with AL to further reduce annotation costs, including MeanTeacher (Tarvainen and Valpola, 2017), VAT (Miyato et al., 2018) and NoTeacher (Unnikrishnan et al., 2021). Huang et al. (2021) combined their proposed COD with MeanTeacher (Tarvainen and Valpola, 2017), demonstrating superior performance. Wang et al. (2022c) combined density-based AL with different existing semi-supervised methods. Results showed that the proposed method outperforms other active learning methods and excels in semi-supervised learning.
Consistency could be also used for sample selection. Gao et al. (2020) introduced a semi-supervised active learning framework. Consistency here was used for both semi-supervised training and evaluating informativeness. In this framework, samples are fed into the model multiple times with random augmentations. The consistency loss of unlabeled samples was implemented by minimizing the variance between multiple outputs. They further selected less consistent samples for annotation. Results showed that combining AL with semi-supervised learning significantly improves performance.
Besides, Zhang et al. (2022a) combined AL with both pseudo-labeling and consistency regularization. The unlabelled images first underwent both strong and weak data augmentations. When the confidence level of the weakly augmented images exceeded a certain threshold, they used these samples for semi-supervised training. Specifically, predictions of the weakly augmented images were assigned as pseudo-labels, and the outputs of the strongly augmented images were forced to be consistent with the pseudo-labels. However, when the confidence level was lower than the threshold, they used these samples for AL. A balanced uncertainty selector and an adversarial instability selector were used to select samples for oracle annotation. They validated the effectiveness of their proposed method in grading metastatic epidural spinal cord compression with MRI images.
@@ 4.2 Self-supervised Learning: Utilizing Pre-trained Model
Integrating semi-supervised learning with AL has achieved successful applications. However, its effectiveness is constrained by the dataset size. This limitation is particularly evident for medical imaging datasets which are relatively small. In clinical practice, plenty of raw medical images are stored in hospital information systems without human annotation. Self-supervised learning (Krishnan et al., 2022) could be a vital tool for mining information hidden in those raw images. Its idea is to train the model with the supervision of the data itself, thus allowing pre-training on a large unlabeled dataset. Many studies have shown self-supervised pre-trained models could achieve impressive performance by finetuning on a few randomly selected labeled samples in medical image analysis (Azizi et al., 2021; Tang et al., 2022b). A natural expectation is to integrate active learning strategies with self-supervised learning, aiming for higher annotation efficiency over mere random sampling (Lüth et al., 2024). Besides, these models could also act as a powerful feature extractor, which provides good initialization for AL. In this section, we will first introduce how self-supervised models solve the cold-start problem in AL and then explore different ways of integrating AL with self-supervised learning.
@@ 4.2.1 Cold-start Problem in Active Learning
Current AL methods usually require an initial labeled dataset to train the model for start and ensure reliable informativeness evaluation. However, when the initial labeled set is small or even absent, the performance of these AL methods drops dramatically, sometimes even worse than random sampling (Chen et al., 2023b; Hacohen et al., 2022; Yehuda et al., 2022). Studies also showed that simply integrating self-supervised learning with AL baselines leads to inferior performance than random sampling (Bengar et al., 2021; Xie et al., 2023a). This phenomenon is known as the cold-start problem which commonly exists in AL of various domains, including medical image analysis (Liu et al., 2023a). Tackling the cold-start problem is vital for improving the efficacy of AL, especially in the medical domains where annotation costs are extremely high. A key solution to the cold-start problem in AL is selecting the optimal set of initial labeled samples, which requires different strategies than the existing AL methods.
@@
Early attempts focused on utilizing the fully supervised pre-trained models to address the cold-start problem in AL. Zhou et al. (2017) and their subsequent work (Zhou et al., 2021c) used ImageNet pre-trained models to select samples for annotation from completely unlabeled datasets in medical image analysis. They combined entropy and disagreement as informativeness metrics, where the disagreement was the KL divergence of prediction probabilities between different patches of the same sample. They also introduced randomness to balance exploration and exploitation. Experiments on two colonoscopy datasets and a CT pulmonary embolism detection dataset showed superior performance than other competitors.
Self-supervised pre-trained models offer a good initialization for effectively tackling the cold-start problem in AL. ALPS (Yuan et al., 2020a) was the first to introduce the cold-start problem in AL and employed self-supervised pre-trained models to address this issue. Based on a contrastive learning feature extractor, CALR (Jin et al., 2022a) employed BIRCH clustering and chose the samples with maximum information density within each cluster for labeling. Compared to k-Means, BIRCH clustering is less sensitive to outliers and can further identify noisy samples. TypiClust (Hacohen et al., 2022) theoretically proved that querying typical samples is more beneficial for a low annotation budget. Therefore, based on self-supervised features, TypiClust selected samples from high-density areas of each k-Means cluster. Beyond that, Yehuda et al. (2022) employed a graph-based greedy algorithm to select the optimal initial samples based on self-supervised features. In CT segmentation, Nath et al. (2022) proposed ProxyRank which designed new pretext tasks for self-supervised pre-training. The model was trained to learn the threshold segmentation by an abdominal soft-tissue window. Results indicated that the proposed method significantly outperforms random sampling in selecting initial samples. To benchmark the effectiveness of different cold-start AL methods in 3D medical image segmentation, Liu et al. (2023a) reproduced ALPS, CALR, TypiClust, and ProxyRank on five MSD datasets (Antonelli et al., 2022). Results showed that TypiClust stands out from the four competitors. However, no method consistently outperformed random selection on all five datasets, which calls for further exploration of cold-start AL in medical image analysis.
@@ 4.2.2 Combination of Active Learning and Self-supervised Learning
Features: The simplest way is leveraging the high-quality features of the self-supervised pre-trained models. Many studies are based on a powerful self-supervised feature extractor (Pourahmadi et al., 2021; Jin et al., 2022a; Hacohen et al., 2022; Yehuda et al., 2022).
@@
Pretext tasks in self-supervised learning are designed to derive supervision directly from the data itself. Solving these pretext tasks on large-scale unlabeled data, the model acquires useful feature representations that reflect data characteristics. Different pretext tasks correspond to different pre-training paradigms, the typical ones including rotation prediction (Gidaris et al., 2018), contrastive learning (He et al., 2020), and masked image modeling (He et al., 2022), etc. Related works generally employed the loss of pretext task for AL. In Chen et al. (2023b), the loss of contrastive learning was used to tackle cold-start problem for AL in medical image analysis. They assumed that samples with higher losses are more representative of the data distribution. Specifically, they pre-trained on the target dataset with momentum contrastive learning (He et al., 2020), and then used k-Means clustering to partition the unlabeled data into multiple clusters, selecting the samples with the highest contrastive loss within each cluster for annotation. They then selected samples with the highest contrastive loss in each cluster for annotation. The proposed method addressed the class imbalance caused by the bias of traditional AL methods, and the failure to detect anomalies when the number of initially labeled datasets was limited. This method showed superior performance in PathMNIST, OrganMNIST, and BloodMNIST (Yang et al., 2023). Yi et al. (2022) found a strong correlation between the loss of pretext tasks and the loss of downstream tasks. Thus, they initially focused on annotating samples with higher loss of pretext tasks and later shifted to those with lower loss. Results showed that rotation prediction performed the best among different pretext tasks.
@@
Others: Furthermore, recent works leverage self-supervised learning in other ways for AL. Zhang et al. (2022b) introduced one-bit annotation into AL for classification tasks. In this setting, oracles only returned whether the prediction was right or wrong rather than its specific class label. Contrastive learning was adopted to pull the correct predictions closer and push away wrong predictions from their predicted classes. Results indicated that the proposed method outperforms other AL methods regarding bit information. Du et al. (2021) integrated contrastive learning into AL to address the class distribution mismatch, where unlabeled data includes samples out of the class distribution of the labeled dataset. In this work, contrastive learning was adopted to filter samples of mismatched classes and highlight sample informativeness by carefully setting negative samples. Their extended work Du et al. (2022) provided more theoretical analysis and experimental results and also integrated existing label information into the proposed framework.